OpenStudy (anonymous):

Given A=16 degrees b=12 and a=10, use the law of sines to solve the triangle if possible for the value of c. If two solutions exist, find both. round answer to two decimal places?

OpenStudy (ranga):

a/sin(A) = b/sin(B) = c/sin(C) ----- (1) a = 10, A = 16 degrees, b = 12 Use a/sin(A) = b/sin(B). Plug in the above three values and solve for B. There may be more than one value for angle B. As long as 0 < B < 180 they are all good. For each B, find angle C (three angles of a triangle add to 180 degrees). Then find the corresponding value of c from equation (1)

OpenStudy (anonymous):

so i got B=19.32 degrees so there can be more than one? how do i check for more?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

sin(180 - 19.32) will give the same value as sin(19.32)